Hello
I have the answers to the following questions but am not sure of the process to solve them. Thanks for the help.
2 people throw a ball at the same time. One of them hits the target 55 out of 100 times, and the other one hits it 95 out of 100 times.
What is the probability that exactly one of the throws Will hit the target?
(answer: 91 / 200 )
What is the probability that at least one of the throws will hit the target?
( 391 / 400 )

Answer by Boreal(15235)

The probability that one only hits is (0.55)(0.05)=0.0275+ (0.45)(0.95)=0.4275 The first term of 0.55*0.05 is the probability the first one hits and the second one doesn't hit. They are independent, so the probability is multiplied. The second is just the opposite.
They add up to 0.4550, or 45.5%, which is 91/200.
At least one is 1-probability that both miss. That is 0.45*0.05=0.0225
That is 2.25% or 2.25 per 100 or 9 per 400 that both miss. Subtract 9 from 400 and 391/400 is the probability that at least 1 will hit.